The physician ordered Ritodrine IV 50 mcg/min. The pharmacy sent up Ritodrine 150 mg premixed in 500mL D5W. You would set your IV pump at ________mL/hour.
Answer #1

ordered dose=50 mcg/min

available dose=150 mg /500 ml D

flow rate=?

conversion:

50 mcg=0.05 mg ( as 1 mcg=0.001 mg)

if 150mg in 500ml

then 0.05 mg in Xml

150*X=500*0.05

X=500*0.05//150=0.1666 ml

0.1666 ml/min to be administered

flow rate=volume in ml/ time in hour

( 1 hour = 60 minutes)

(therefore 1 min=0.0166 hrs)

flow rate=0.1666/0.0166=1.003 ml/hour Ritodrine to be given.